OpenStudy (anonymous):

Simplify the expression so that it uses positive exponents, only. (x^-1/2y^2)(x^5/2y^-4)

OpenStudy (anonymous):

its (1/x/2y^2)(x/5/2/y^4 equals (1/x*2y^2)(x/10y^4) so its \[(\frac{ 1 }{ x2y^2 })(\frac{ x }{ 10y^4 })\]

OpenStudy (anonymous):

can u simplify it now?

OpenStudy (anonymous):

if you multiply it you get: \[\frac{ 1 }{ 20y^6 }\]

OpenStudy (anonymous):

so the answer is: \[\frac{ 1 }{ 20y^6 }\]